\begin{thebibliography}{10}

\bibitem{ALDROUBISUNWAI}
Akram Aldroubi, Qiyu Sun, and Wai-Shing Tang.
\newblock {$p$}-frames and shift invariant subspaces of {$L^p$}.
\newblock {\em J. Fourier Anal. Appl.}, 7(1):1--21, 2001.

\bibitem{BALANCASAZZAHEILLANDAU}
Radu Balan, Peter~G. Casazza, Christopher Heil, and Zeph Landau.
\newblock Density, overcompleteness, and localization of frames. {I}. {T}heory.
\newblock {\em J. Fourier Anal. Appl.}, 12(2):105--143, 2006.

\bibitem{BARANOVDYAKONOV}
Anton Baranov and Konstantin Dyakonov.
\newblock The {F}eichtinger conjecture for reproducing kernels in model
  subspaces.
\newblock {\em J. Geom. Anal.}, 21(2):276--287, 2011.

\bibitem{BENEDETTOWALNUT}
John~J. Benedetto and David~F. Walnut.
\newblock Gabor frames for {$L^2$} and related spaces.
\newblock In {\em Wavelets: mathematics and applications}, Stud. Adv. Math.,
  pages 97--162. CRC, Boca Raton, FL, 1994.

\bibitem{BOWNIK}
Marcin Bownik.
\newblock The {K}adison-{S}inger problem.
\newblock In {\em Frames and harmonic analysis}, volume 706 of {\em Contemp.
  Math.}, pages 63--92. Amer. Math. Soc., [Providence], RI, 2018.

\bibitem{BOWNIKSPEEGLE}
Marcin Bownik and Darrin Speegle.
\newblock The {F}eichtinger conjecture for wavelet frames, {G}abor frames and
  frames of translates.
\newblock {\em Canad. J. Math.}, 58(6):1121--1143, 2006.

\bibitem{CASAZZADILWORTHODELL}
P.~G. Casazza, S.~J. Dilworth, E.~Odell, Th. Schlumprecht, and A.~Zs\'{a}k.
\newblock Coefficient quantization for frames in {B}anach spaces.
\newblock {\em J. Math. Anal. Appl.}, 348(1):66--86, 2008.

\bibitem{CASAZZACHRISTENSENLINDNERVERSHYNIN}
Peter~G. Casazza, Ole Christensen, Alexander~M. Lindner, and Roman Vershynin.
\newblock Frames and the {F}eichtinger conjecture.
\newblock {\em Proc. Amer. Math. Soc.}, 133(4):1025--1033, 2005.

\bibitem{CASAZZAEDIDIN}
Peter~G. Casazza and Dan Edidin.
\newblock Equivalents of the {K}adison-{S}inger problem.
\newblock In {\em Function spaces}, volume 435 of {\em Contemp. Math.}, pages
  123--142. Amer. Math. Soc., Providence, RI, 2007.

\bibitem{CASAZZAFICKUSTREMAINWEBER}
Peter~G. Casazza, Matthew Fickus, Janet~C. Tremain, and Eric Weber.
\newblock The {K}adison-{S}inger problem in mathematics and engineering: a
  detailed account.
\newblock In {\em Operator theory, operator algebras, and applications}, volume
  414 of {\em Contemp. Math.}, pages 299--355. Amer. Math. Soc., Providence,
  RI, 2006.

\bibitem{CASAZZAHANLARSON}
Peter~G. Casazza, Deguang Han, and David~R. Larson.
\newblock Frames for {B}anach spaces.
\newblock In {\em The functional and harmonic analysis of wavelets and frames
  ({S}an {A}ntonio, {TX}, 1999)}, volume 247 of {\em Contemp. Math.}, pages
  149--182. Amer. Math. Soc., Providence, RI, 1999.

\bibitem{CASAZZAKUTYNIOKSPEEGLETREMAIN}
Peter~G. Casazza, Gitta Kutyniok, Darrin Speegle, and Janet~C. Tremain.
\newblock A decomposition theorem for frames and the {F}eichtinger conjecture.
\newblock {\em Proc. Amer. Math. Soc.}, 136(6):2043--2053, 2008.

\bibitem{CHRISTENSEN2014}
Ole Christensen.
\newblock Six (seven) problems in frame theory.
\newblock In {\em New perspectives on approximation and sampling theory}, Appl.
  Numer. Harmon. Anal., pages 337--358. Birkh\"{a}user/Springer, Cham, 2014.

\bibitem{CHRISTENSEN}
Ole Christensen.
\newblock {\em An introduction to frames and {R}iesz bases}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user/Springer,
  second edition, 2016.

\bibitem{CHRISTENSENSTOEVA}
Ole Christensen and Diana~T. Stoeva.
\newblock {$p$}-frames in separable {B}anach spaces.
\newblock {\em Adv. Comput. Math.}, 18(2-4):117--126, 2003.

\bibitem{DUFFINSCHAEFFER}
R.~J. Duffin and A.~C. Schaeffer.
\newblock A class of nonharmonic {F}ourier series.
\newblock {\em Trans. Amer. Math. Soc.}, 72:341--366, 1952.

\bibitem{FEICHTINGERGROCHENIG}
Hans~G. Feichtinger and K.~Gr\"{o}chenig.
\newblock Gabor frames and time-frequency analysis of distributions.
\newblock {\em J. Funct. Anal.}, 146(2):464--495, 1997.

\bibitem{FEICHTINGERSTROHMER2}
Hans~G. Feichtinger and Thomas Strohmer, editors.
\newblock {\em Gabor analysis and algorithms: Theory and application}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user Boston, Inc.,
  Boston, MA, 1998.

\bibitem{FEICHTINGERSTROHMER}
Hans~G. Feichtinger and Thomas Strohmer, editors.
\newblock {\em Advances in {G}abor analysis}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user Boston, Inc.,
  Boston, MA, 2003.

\bibitem{FRNASIERGROCHENIG}
Massimo Fornasier and Karlheinz Gr\"{o}chenig.
\newblock Intrinsic localization of frames.
\newblock {\em Constr. Approx.}, 22(3):395--415, 2005.

\bibitem{FREEMANODELLSCHIUMPRECHT}
D.~Freeman, E.~Odell, Th. Schlumprecht, and A.~Zs\'{a}k.
\newblock Unconditional structures of translates for {$L_p(\mathbb{R}^d)$}.
\newblock {\em Israel J. Math.}, 203(1):189--209, 2014.

\bibitem{GROCHENIGBOOK}
Karlheinz Gr\"{o}chenig.
\newblock {\em Foundations of time-frequency analysis}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user Boston, Inc.,
  Boston, MA, 2001.

\bibitem{GROCHENIG}
Karlheinz Gr\"{o}chenig.
\newblock Localized frames are finite unions of {R}iesz sequences.
\newblock {\em Adv. Comput. Math.}, 18(2-4):149--157, 2003.

\bibitem{GROCHENIG2004}
Karlheinz Gr\"{o}chenig.
\newblock Localization of frames, {B}anach frames, and the invertibility of the
  frame operator.
\newblock {\em J. Fourier Anal. Appl.}, 10(2):105--132, 2004.

\bibitem{GROCHENIG2007}
Karlheinz Gr\"{o}chenig.
\newblock Gabor frames without inequalities.
\newblock {\em Int. Math. Res. Not. IMRN}, (23):Art. ID rnm111, 21, 2007.

\bibitem{GROCHENIG2014}
Karlheinz Gr\"{o}chenig.
\newblock The mystery of {G}abor frames.
\newblock {\em J. Fourier Anal. Appl.}, 20(4):865--895, 2014.

\bibitem{GROCHENIG2015}
Karlheinz Gr\"{o}chenig, Joaquim Ortega-Cerd\`a, and Jos\'{e}~Luis Romero.
\newblock Deformation of {G}abor systems.
\newblock {\em Adv. Math.}, 277:388--425, 2015.

\bibitem{HEIL}
Christopher Heil.
\newblock History and evolution of the density theorem for {G}abor frames.
\newblock {\em J. Fourier Anal. Appl.}, 13(2):113--166, 2007.

\bibitem{LATAPAULSEN}
Sneh Lata and Vern Paulsen.
\newblock The {F}eichtinger conjecture and reproducing kernel {H}ilbert spaces.
\newblock {\em Indiana Univ. Math. J.}, 60(4):1303--1317, 2011.

\bibitem{LAWTON}
Wayne Lawton.
\newblock The {F}eichtinger conjecture for exponentials.
\newblock {\em J. Nonlinear Anal. Optim.}, 2(1):131--140, 2011.

\bibitem{KRISHNA}
K.~Mahesh~Krishna.
\newblock Feichtinger conjectures, ${R}_\varepsilon$-conjectures and {W}eaver's
  conjectures for {B}anach spaces.
\newblock {\em arXiv.org/2201.00125v1 [math.FA] 1 January}, 2022.

\bibitem{KRISHNAJOHNSON2}
K.~Mahesh~Krishna and P.~Sam Johnson.
\newblock Dilation theorem for p-approximate {S}chauder frames for separable
  {B}anach spaces.
\newblock {\em Palest. J. Math.}, 11(2):384--394, 2022.

\bibitem{MARCUSSRIVASTAVA}
Adam Marcus and Nikhil Srivastava.
\newblock The solution of the {K}adison-{S}inger problem.
\newblock In {\em Current developments in mathematics 2016}, pages 111--143.
  Int. Press, Somerville, MA, 2018.

\bibitem{MARCUSSPIELMANSRIVASTAVA}
Adam~W. Marcus, Daniel~A. Spielman, and Nikhil Srivastava.
\newblock Interlacing families {II}: {M}ixed characteristic polynomials and the
  {K}adison-{S}inger problem.
\newblock {\em Ann. of Math. (2)}, 182(1):327--350, 2015.

\bibitem{SPEEGLE2008}
Darrin Speegle.
\newblock Uniform partitions of frames of exponentials into {R}iesz sequences.
\newblock {\em J. Math. Anal. Appl.}, 348(2):739--745, 2008.

\bibitem{STOEVA}
Diana~T. Stoeva.
\newblock On a characterization of {R}iesz bases via biorthogonal sequences.
\newblock {\em J. Fourier Anal. Appl.}, 26(4):Paper No. 67, 5, 2020.

\bibitem{TIMOTIN}
Dan Timotin.
\newblock The solution of the {K}adison-{S}inger problem.
\newblock In {\em Recent advances in operator theory and operator algebras},
  pages 117--148. CRC Press, Boca Raton, FL, 2018.

\bibitem{WEAVER}
Nik Weaver.
\newblock The {K}adison-{S}inger problem in discrepancy theory.
\newblock {\em Discrete Math.}, 278(1-3):227--239, 2004.

\bibitem{WEBER}
Eric Weber.
\newblock Algebraic aspects of the paving and {F}eichtinger conjectures.
\newblock In {\em Topics in operator theory. {V}olume 1. {O}perators, matrices
  and analytic functions}, volume 202 of {\em Oper. Theory Adv. Appl.}, pages
  569--578. Birkh\"{a}user Verlag, Basel, 2010.

\end{thebibliography}

\documentclass{amsart}
\usepackage{setspace}
\usepackage{a4}
\usepackage{amsthm,latexsym}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{cite}
\usepackage{enumerate}
\usepackage[mathscr]{euscript}
\usepackage{mathtools}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{conjecture}[theorem]{Conjecture}

\newtheorem{definition}[theorem]{Definition}

\title{This is the title}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{mathtools}
\usetikzlibrary{cd}
\usepackage{fancyhdr}

\begin{document}
	\hrule\hrule\hrule\hrule\hrule
	\vspace{0.3cm}	
	\begin{center}
		{\bf{LOCALIZED BOUNDED BELOW APPROXIMATE SCHAUDER FRAMES ARE FINITE UNIONS OF APPROXIMATE RIESZ SEQUENCES}}\\
		\vspace{0.3cm}
		\hrule\hrule\hrule\hrule\hrule
		\vspace{0.3cm}
		\textbf{K. MAHESH KRISHNA}\\
		Post Doctoral Fellow \\
		Statistics and Mathematics Unit\\
		Indian Statistical Institute, Bangalore Centre\\
		Karnataka 560 059, India\\
		Email: kmaheshak@gmail.com\\
		
		Date: \today
	\end{center}


\hrule
\vspace{0.5cm}
\textbf{Abstract}: Based on the truth of Feichtinger conjecture by Marcus, Spielman and Srivastava \textit{[Ann. of Math. (2), 2015]} and from the localized version by Gr\"{o}chenig  \textit{[Adv. Comput. Math., 2003]}, we introduce the notion of localization of approximate Schauder frames (ASFs) and approximate Riesz sequences (ARSs). We show that localized bounded below ASFs are finite unions of ARSs.

\textbf{Keywords}:  Feichtinger conjecture, Frame, Riesz sequence, Localization.

\textbf{Mathematics Subject Classification (2020)}: 42C15, 46A45, 46B45.
\vspace{0.5cm}
\hrule 

\section{Introduction}
In the beginning years of $21^{th}$ century, Prof. Feichtinger  formulated the following conjecture based on his extensive work on Gabor/Weyl-Heisenberg frames (see \cite{CHRISTENSEN2014} for the history and  \cite{STOEVA, DUFFINSCHAEFFER, CHRISTENSEN, FEICHTINGERSTROHMER, FEICHTINGERSTROHMER2, HEIL, FEICHTINGERGROCHENIG, GROCHENIG2007, GROCHENIG2014, BENEDETTOWALNUT, GROCHENIGBOOK, GROCHENIG2015} for general theory).
\begin{conjecture}\cite{CASAZZACHRISTENSENLINDNERVERSHYNIN} \textbf{(Feichtinger Conjecture/Marcus-Spielman-Srivastava Theorem)}\label{FCON}
	\textbf{Let  $\{\tau_n\}_n$ be  a frame for  a Hilbert space  $\mathcal{H}$ such that 
		\begin{align*}
			0<\inf_{n\in \mathbb{N}}\|\tau_n\|.
		\end{align*}
		Then $\{\tau_n\}_n$ can be partitioned into a finite union of Riesz sequences for $\mathcal{H}$.}
\end{conjecture}
First breakthrough which supported  Conjecture \ref{FCON} occurred when Gr\"{o}chenig  proved it for intrinsically  localized frames \cite{GROCHENIG}. Shortly afterwords, it has been verified for certain classes of $\ell^1$-self-localized frames, wavelet frames, Gabor frames, frames of translates, frames formed by reproducing kernels  and exponential frames/frames of exponentials \cite{BOWNIKSPEEGLE, BALANCASAZZAHEILLANDAU, SPEEGLE2008, LATAPAULSEN, BARANOVDYAKONOV, LAWTON, WEBER}.  Conjecture \ref{FCON} received great attention after establishing its   equivalence  with Kadison-Singer conjecture \cite{CASAZZAFICKUSTREMAINWEBER, CASAZZAEDIDIN, CASAZZAKUTYNIOKSPEEGLETREMAIN}. Finally, the Feichtinger conjecture has been solved fully by  resolving Weaver's conjecture by Marcus, Spielman, and Srivastava in 2013 \cite{MARCUSSPIELMANSRIVASTAVA, WEAVER, BOWNIK, TIMOTIN, MARCUSSRIVASTAVA}. In this paper, we formulate a Banach space version of Conjecture \ref{FCON} and prove it for bounded below intrinsically localized ASFs (Theorem \ref{THM}).

\section{Localized bounded below ASFs are finite unions of  ARSs}
We consider the following most general notion of  approximate Schauder frames. In the entire paper,  $\mathcal{X}$ is a  separable Banach space and $\mathcal{X}^*$ is its dual. 

\begin{definition}\cite{CASAZZAHANLARSON, CASAZZADILWORTHODELL, FREEMANODELLSCHIUMPRECHT}
	Let $ \{\tau_n \}_{n}$ be a collection in $\mathcal{X}$ and $ \{f_n \}_{n}$ be a collection in $\mathcal{X}^*$.	The pair  $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $  	is said to be an \textbf{approximate Schauder frame} (we write ASF)  for $\mathcal{X}$ if 
the \textbf{frame operator} 
\begin{align*}
S_{f, \tau}:\mathcal{X}\ni x \mapsto S_{f, \tau}x\coloneqq \sum_{n=1}^\infty f_n(x)\tau_n \in\mathcal{X}	
\end{align*}
 is a well-defined bounded linear invertible operator. 
		
\end{definition}
We use the following notion of `bounded below' for ASFs.
\begin{definition}
	An ASF $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $  for $\mathcal{X}$ is said to be \textbf{bounded below} if 
	\begin{align*}
		\inf_{n \in \mathbb{N}}|f_n(\tau_n)|>0.
	\end{align*}
\end{definition}
Motivated from the definition of localization of frames \cite{GROCHENIG2004, FRNASIERGROCHENIG}, we introduce the following notion.
\begin{definition}
An ASF $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $  for $\mathcal{X}$ is said to be \textbf{intrinsically/self localized} if 	there exist $s>1$ and $A>0$ such that 
\begin{align*}
|f_n(\tau_m)|\leq \frac{A}{(1+|n-m|)^s}, \quad \forall n, m \in \mathbb{N}.
\end{align*}
\end{definition}
Two notions of Riesz sequences for Banach spaces exist in literature, see 
\cite{CHRISTENSENSTOEVA, ALDROUBISUNWAI} and \cite{KRISHNAJOHNSON2}. Here we define another.
\begin{definition}\label{SRB}
An ASF $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $  for $\mathcal{X}$ is  said to be an \textbf{approximate Riesz sequence} (we write ARS) if there exists a finite partition $Q_1, \dots, Q_N$ of $\mathbb{N}$ such that 
\begin{align*}
	\mathbb{N}=\bigcup_{j=1}^NQ_j
\end{align*}
and for each $1\leq j \leq N$, 
\begin{align*}
	\inf_{n \in Q_j}\left(|f_n(\tau_n)|-\sum_{m \in Q_j, m \neq n}|f_n(\tau_m)|\right)>0.
\end{align*}
\end{definition}
Note that for Hilbert spaces, if $\{f_n \}_{n}$ is determined by $\{\tau_n \}_{n}$ (Riesz representation), then Definition \ref{SRB} is equivalent (due to positivity) to the definition of Riesz sequence (see\cite{GROCHENIG}). We now formulate the following conjecture (some other are formulated in \cite{KRISHNA}).

\begin{conjecture}\label{CONJECTURE}
	\textbf{Every bounded below ASF can be partitioned  as a finite union of ARBs.}
\end{conjecture}
We now prove Conjecture \ref{CONJECTURE} for intrinsically localized ASFs with the help of following result.
\begin{theorem}\cite{GROCHENIG}\label{GT}
	\begin{enumerate}[\upshape(i)]
		\item For every $s>1$, 
		\begin{align*}
			D_s\coloneqq 	\sup_{x \in \mathbb{R}} \sum_{n=1} ^\infty \frac{1}{(1+|n-x|)^s}<\infty.
		\end{align*}
	\item  For every $s>1$, there exists a $C_s>0$ (which does not depend on  $\delta$) such that 	
	\begin{align*}
		\sup_{m \in \mathbb{N}} \sum_{n \in \mathbb{N}, n \neq m} \frac{1}{(1+|n-m|)^s}\leq \frac{C_s}{\delta^s}, 
	\end{align*}
whenever  
\begin{align*}
	\inf_{n, m \in \mathbb{N}, n \neq m}|n-m|\geq \delta.
\end{align*}
	\end{enumerate}
\end{theorem}
\begin{theorem}\label{THM}
Conjecture \ref{CONJECTURE} holds for intrinsically localized bounded below ASFs.	
\end{theorem}
\begin{proof}
Our proof is highly motivated from \cite{GROCHENIG}. Let $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $  be an intrinsically localized bounded below ASF for  $\mathcal{X}$.  
Define 
\begin{align*}
C\coloneqq \inf_{n \in \mathbb{N}}|f_n(\tau_n)|>0.
\end{align*}	
Since $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $  intrinsically localized	there exist $s>1$ and $A>0$ such that 
\begin{align*}
	|f_n(\tau_m)|\leq \frac{A}{(1+|n-m|)^s}, \quad \forall n, m \in \mathbb{N}.
\end{align*}
Let $C_s$ be the constant  as in Theorem \ref{GT}. Choose natural number $M$ such that 
\begin{align*}
	\frac{AC_s}{M^s}\leq \frac{C}{2}.
\end{align*}
We now partition $\mathbb{N}$ into $Q_1, \dots, Q_N$ such that 
\begin{align*}
	\mathbb{N}\coloneqq \bigcup_{j=1}^NQ_j
\end{align*}
and for each $ 1\leq j\leq N$,
\begin{align*}
	\inf_{n, m \in Q_j, n \neq m}|n-m|\geq M.	
\end{align*}
(Note that there are infinitely many partitions of $\mathbb{N}$ of this type.) Let $ 1\leq j\leq N$. Then using (ii) in Theorem \ref{GT}
\begin{align*}
	\sup_{n \in Q_j}\sum_{m \in Q_j, m \neq n}|f_n(\tau_m)|\leq A \sup_{n \in Q_j}\sum_{m \in Q_j, m \neq n}\frac{1}{(1+|n-m|)^s}\leq A \frac{C_s}{M^s}\leq \frac{C}{2}.
\end{align*}
Therefore for each fixed $ 1\leq j\leq N$
\begin{align*}
\inf_{n \in Q_j}\left(|f_n(\tau_n)|-\sum_{m \in Q_j, m \neq n}|f_n(\tau_m)|\right)&=\inf_{n \in Q_j}|f_n(\tau_n)|-\sup_{n \in Q_j}\sum_{m \in Q_j, m \neq n}|f_n(\tau_m)|\\
&\geq C-\frac{C}{2}=\frac{C}{2}>0.
\end{align*}
Since $j$ was arbitrary, we get the theorem.
\end{proof}

  \bibliographystyle{plain}
 \bibliography{reference.bib}

\end{document}